# Semi-empirical constraints on the HI mass function of star-forming galaxies and $\Omega_{\rm HI}$ at $z\sim 0.37$ from interferometric surveys

### v1

@@ 5.3 Discussion
As already anticipated, the main findings of this work point towards an evolution in the HIMF and in the $\Omega_{\rm HI}$ parameters with respect to the fiducial ALFALFA HIMF at $z\sim 0$. This fact is consistently inherited from the scaling relation from Bianchetti et al. (2025), which also features a clear evolution with respect to literature results at $z\sim 0$. In this picture, and including the HIMF from Chowdhury et al. (2024), the H I content of galaxies would evolve gradually and smoothly from $z\sim 0$ to $z\sim 1$. This scenario is in tension with the results from Bera et al. (2022) and from Paul et al. (2023). In fact, both references promote a HIMF with similar normalization at $\log_{10}(M_{\rm HI}/{\rm M}_{\odot})\sim 9-9.5$ as from this work, but with a quicker fall-off towards high $M_{\rm HI}$. This typically results in a lower value for $\Omega_{\rm HI}$. In the case of Bera et al. (2022), the results is obtained with a similar approach as ours, but the authors therein used a different scaling relation to constrain to pass from $M_{B}$ to $M_{\rm HI}$. In addition, the sample size is a factor $\sim 10$ smaller than the one used in Bianchetti et al. (2025), so cosmic variance could be affecting at least partially the results when measuring H I masses from stacking. Eventually, the selection of star-forming galaxies performed by Bera et al. (2022) and Bianchetti et al. (2025) are based on different criteria based on colors. These aspects altogether may explain the observed differences, although we regard this comparison as not conclusive as in general more investigation with larger sample sizes is required. With respect to Paul et al. (2023), the HIMF obtained in that work are based on H I intensity mapping, i.e. a radically different approach from the one adopted herein. Therefore, it is difficult to establish a fair comparison in this case. Nonetheless, it is worth mentioning that the result obtained via H I intensity mapping is in good agreement with the HIMF obtained by Bera et al. (2022).
@@
We notice that our results have a non-negligible dependence on the model adopted to describe the $M_{\rm HI}-M_{\star}$ relation. In fact, the linear model tends to produce HIMF with higher normalization and larger values for $\Omega_{\rm HI}$ than the polynomial model. This is not surprising since, as discussed, the polynomial model steepens towards low $M_{\star}$ where a greater number of galaxies is found. In addition, a linear model extrapolates the $M_{\rm HI}-M_{\star}$ relation to arbitrarily high value of $M_{\rm HI}$. However, results at $z\sim 0$ show that massive galaxies feature a flattening in $M_{\rm HI}$ at the high-mass end of the $M_{\rm HI}-M_{\star}$ relation, due to a plethora of phenomena shaping galaxy evolution, mainly gas depletion via star formation. In this work, we find evidence for the polynomial model to better fit the data points from stacking. In particular, the best-fitting second-order polynomial model for the $M_{\rm HI}$ relation is found to have second-order coefficient $a<0$ at a significance $>10\sigma$ in all the studied cases. While we regard this finding as non-conclusive given the limited amount of data points available for the fit, the trend suggested by the best-fitting polynomial fit qualitatively reproduces the one described by a double power law at $z\sim 0$ (Maddox et al., 2015; Parkash et al., 2018; Pan et al., 2023). In particular, the polynomial model with $a<0$ accounts for the high-mass flattening of the $M_{\rm HI}-M_{\star}$ scaling relation. For all these reasons we regard the polynomial model as the fiducial model for this work.
@@ 5.4 Comparison with simulations
All the other feedback models converge well within each other and are in well agreement with our results at $\log_{10}(M_{\rm HI}/{\rm M_{\odot}})\sim 9.7$, but tend to diverge towards high $M_{\rm HI}$ values. The noX model underpredicts our results, while the noAGN is in good agreement with them, and the noJet mode overpredicts the HIMF, respectively. This indicates that, as anticipated, the specific AGN feedback model is a major driver of the massive end of the HIMF. In particular, it is interesting to notice that the noAGN model reproduces very well the HIMF from this work. The finding suggests a scenario in which neglecting the feedback from supermassive black holes allows to better model the HIMF at the probed H I masses. This is consistent with the results from Gensior et al. (2023, 2024), where the authors considered tens of Milky-Way-like galaxies from the EMP-Pathfinder (Reina-Campos et al., 2022) and FIREbox (Feldmann et al., 2023) cosmological simulations and showed that they reproduce well the observed H I properties of nearby galaxies without the need of AGN feedback. Therefore, despite the large uncertainty associated to feedback models in current cosmological hydrodynamic simulations, we argue that a detailed study of these feedback mechanisms is crucial to properly understand which physical mechanism shapes the HIMF at the different H I masses.
@@ 6 Summary and conclusions
In this work we report a novel indirect semi-empirical estimation of the HIMF and of the H I the cosmic density parameters $\Omega_{\rm HI}$ for star-forming galaxies at $z\sim 0.37$. In particular, we combine the most recent measurements of the stellar mass function of star-forming galaxies in the COSMOS field at $0.50.2<z<$ 0.5 (Weaver et al., 2023) and of the $M_{\rm HI}-M_{\star}$ scaling relation of star-forming galaxies at $0.490.22<z<$ 0.49 (median redshift $z\sim 0.37$) by Bianchetti et al. (2025) from the combined H I stacking analysis of two state-of-the-art deep H I interferometric surveys MIGHTEE (Jarvis et al., 2016) and CHILES (Fernández et al., 2013), both covering the COSMOS field. In particular, we adopt a Monte Carlo approach following the one pioneered by Bera et al. (2022) consisting of drawing samples of the stellar mass function, generate a synthetic stellar mass sample from each stellar mass function, and convert $M_{\star}$ to $M_{\rm HI}$ by assuming a $M_{\rm HI}-M_{\star}$ scaling relation with an adequate scatter. We compute the median and the standard deviation of the sample of HIMF, which defines our ‘measured’ HIMF. In addition, for each H I mass sample, we build the corresponding HIMF and fit a Schechter profile via a Markov Chain Monte Carlo procedure using the emcee affine-invariant sampler (Foreman-Mackey et al., 2013). This allows us to determine the full posterior distributions of the best-fitting Schechter parameters and hence, to determine not only the median values but also their associated errors after having adequately propagated all the uncertainties in our forward modelling.
@@
  • we model the $M_{\rm HI}-M_{\star}$ scaling relation by fitting the data points from Bianchetti et al. (2025) using both a linear and a second-order polynomial model, finding a preference for the second. In particular, the second-order coefficient $a$ is found to be negative $a<0$ at a significance $>10\sigma$. This fact has important implications for galaxy evolution. In fact, this suggests that the $M_{\rm HI}-M_{\star}$ should decrease more rapidly than a linear model towards low $M_{\star}$, and flatten towards high $M_{\star}$;
  • we find a global evolution in the normalization of the HIMF from the $z\sim 0$ to $z\sim 1$. In particular, the resulting HIMF from this work has larger normalization than the one reported by ALFALFA at $z\sim 0$, and smaller normalization than the one reported by Chowdhury et al. (2024) at $z\sim 1$;
@@
  • the resulting values for $\Omega_{\rm HI}$ obtained by assuming the linear and the polynomial model for the scaling relation differ by $\sim 1.1\sigma$ in the combined stacking case.
@@ Appendix A Consistency of the HIMF from MIGHTEE and CHILES
Survey $\log_{10}(\tilde{M})$ $\alpha$ $\log_{10}(\tilde{\Phi})$ $\Omega_{\rm HI}\times 10^{-4}$
@@
MIGHTEE $10.13^{+0.08}_{-0.09}$ $-1.38^{+0.11}_{-0.10}$ $-2.24^{+0.13}_{-0.13}$ $8.16^{+0.59}_{-0.54}$
CHILES $10.01^{+0.16}_{-0.15}$ $-1.31^{+0.23}_{-0.20}$ $-2.18^{+0.21}_{-0.25}$ $6.65^{+0.99}_{-0.69}$
Combined $10.16^{+0.08}_{-0.08}$ $-1.48^{+0.10}_{-0.09}$ $-2.41^{+0.12}_{-0.13}$ $7.02^{+0.59}_{-0.52}$
@@
MIGHTEE $9.94^{+0.09}_{-0.09}$ $-1.09^{+0.15}_{-0.14}$ $-1.85^{+0.11}_{-0.13}$ $9.66^{+0.57}_{-0.56}$
CHILES $9.92^{+0.5}_{-0.14}$ $-1.22^{+0.23}_{-0.22}$ $-1.95^{+0.19}_{-0.23}$ $8.06^{+0.95}_{-0.85}$
Combined $10.07^{+0.08}_{-0.08}$ $-1.42^{+0.12}_{-0.11}$ $-2.19^{+0.13}_{-0.14}$ $8.51^{+0.69}_{-0.56}$
@@
The best-fitting values for the parameters of the Schechter model are always in agreement within $1\sigma$ in the combined, MIGHTEE, and CHILES cases. When comparing the results from the linear and the polynomial model, we report a systematic shift of $\log_{10}(\tilde{M})$ towards larger values and of $\alpha$ and $\log_{10}(\tilde{\Phi})$ towards smaller values when passing from the linear to the polynomial model. While these shifts are not statistically significant and are often $<1\sigma$, we observe them systematically for all the three cases. In addition, we show in Appendix F the posterior distributions for the different cases. Figs. 10 and 11 show the resulting posterior distributions for the model parameters of the Schechter profile for the linear and the polynomial model, respectively. The distributions have qualitatively similar shapes, and the ones for CHILES are manifestly broader than the posteriors of the MIGHTEE and the combined cases, as expected from the smaller statistics. We will have a closer look to the effect of the sample size in Appendix E. Figs. 12, 13, and 14 display instead a comparison between the posterior distributions from the linear (blue solid) and the polynomial model (orange dashed), for the combined, MIGHTEE, and CHILES cases, respectively. Here, we clearly appreciate the systematic shift of the posteriors when passing from the linear to the polynomial model.
@@ Appendix B The impact of scatter
Survey $\log_{10}(\tilde{M})$ $\alpha$ $\log_{10}(\tilde{\Phi})$ $\Omega_{\rm HI}\times 10^{-4}$
@@
MIGHTEE $10.13^{+0.08}_{-0.09}$ $-1.38^{+0.11}_{-0.10}$ $-2.24^{+0.13}_{-0.13}$ $8.16^{+0.59}_{-0.54}$
CHILES $10.01^{+0.16}_{-0.15}$ $-1.31^{+0.23}_{-0.20}$ $-2.18^{+0.21}_{-0.25}$ $6.65^{+0.99}_{-0.69}$
Combined $10.16^{+0.08}_{-0.08}$ $-1.48^{+0.10}_{-0.09}$ $-2.41^{+0.12}_{-0.13}$ $7.02^{+0.59}_{-0.52}$
@@
MIGHTEE $10.11^{+0.09}_{-0.10}$ $-1.37^{+0.11}_{-0.10}$ $-2.23^{+0.12}_{-0.14}$ $8.09^{+0.60}_{-0.54}$
CHILES $10.02^{+0.16}_{-0.16}$ $-1.33^{+0.22}_{-0.19}$ $-2.20^{+0.21}_{-0.24}$ $6.67^{+0.98}_{-0.80}$
Combined $10.15^{+0.08}_{-0.08}$ $-1.48^{+0.10}_{-0.09}$ $-2.41^{+0.13}_{-0.13}$ $6.99^{+0.61}_{-0.51}$
@@
Figure 7: Results of testing the impact of an asymmetric scatter for the $M_{\rm HI}-M_{\star}$ relation. Top: $M_{\rm HI}$ distributions (to which we have subtracted the median) in different stellar mass bins: $9.18.9<\log_{10}(M_{\star}/{\rm M_{\odot}})<$ 9.1 (green solid), $9.69.4<\log_{10}(M_{\star}/{\rm M_{\odot}})<$ 9.6 (orange dashed), $10.19.9<\log_{10}(M_{\star}/{\rm M_{\odot}})<$ 10.1 (blue dotted-dashed), $10.510.4<\log_{10}(M_{\star}/{\rm M_{\odot}})<$ 10.5 (purple dotted). We display as a gray dashed a lognormal random distribution obtained by assuming $\sigma=0.45$, which is found to approximate well all the $M_{\rm HI}$ distributions.
@@ Appendix D Completeness effects on the HIMF
Fig. 8 shows the resulting HIMF from assuming the combined scaling relation and the linear (yellow stars) and the polynomial (blue squares) models. The HIMF presents evidence of incompleteness at $\log_{10}(M_{\rm HI}/{\rm M}_{\odot})\lesssim 9$. As stated in Sect. 4.3, we conservatively fit a Schechter profile to the resulting HIMF at $\log_{10}(M_{\rm HI}/{\rm M}_{\odot})>9.2$. The $M_{\rm HI}$ lower limit for the fit $\log_{10}(M_{\rm HI}/{\rm M}_{\odot})=9.2$ is marked by a vertical black dashed line. We display the best-fitting Schechter models obtained adopting the fitting strategy outlined above as solid yellow (linear model) and solid blue (polynomial model lines.
Figure 8: Completeness effects on the HIMF. We display the resulting HIMF obtained by assuming the linear (yellow stars) and the polynomial (blue squares) models for the $M_{\rm HI}-M_{\star}$ relation. The vertical black dashed line stands for the conservative completeness limit $\log_{10}(M_{\rm HI}/{\rm M}_{\odot})=9.2$, and represents the lower limit used for the fit of the Schechter model to the data. We show the best-fitting Schechter models obtained adopting the fitting strategy outlined above as solid yellow (linear model) and solid blue (polynomial model lines. We also display the HIMF from ALFALFA at $z\sim 0$ (green dashed) for comparison.

### v2

@@ 5.3 Discussion
As already anticipated, the main findings of this work point towards a difference in the HIMF and in the $\Omega_{\rm HI}$ parameters with respect to the fiducial ALFALFA HIMF at $z\sim 0$. This fact is consistently inherited from the scaling relation from Bianchetti et al. (2025), which also features a clear deviation with respect to literature results at $z\sim 0$. Assuming that the detected variations are due to actual evolution, and including the HIMF from Chowdhury et al. (2024) in the picture, the H I content of galaxies would evolve gradually and smoothly from $z\sim 0$ to $z\sim 1$. This scenario is in tension with the results from Bera et al. (2022) and from Paul et al. (2023). In fact, both references promote a HIMF with similar normalization at $\log_{10}(M_{\rm HI}/{\rm M}_{\odot})\sim 9-9.5$ as from this work, but with a quicker fall-off towards high $M_{\rm HI}$. This typically results in a lower value for $\Omega_{\rm HI}$. In the case of Bera et al. (2022), the results is obtained with a similar approach as ours, but the authors therein used a different scaling relation to pass from $M_{B}$ to $M_{\rm HI}$. In addition, the sample size is a factor $\sim 10$ smaller than the one used in Bianchetti et al. (2025), so cosmic variance could be affecting at least partially the results when measuring H I masses from stacking. Eventually, the selection of star-forming galaxies performed by Bera et al. (2022) and Bianchetti et al. (2025) are based on different criteria based on colors. These aspects altogether may explain the observed differences, although we regard this comparison as not conclusive as in general more investigation with larger sample sizes is required. With respect to Paul et al. (2023), the HIMF obtained in that work are based on H I intensity mapping, i.e. a radically different approach from the one adopted herein. Therefore, it is difficult to establish a fair comparison in this case. Nonetheless, it is worth mentioning that the result obtained via H I intensity mapping is in good agreement with the HIMF obtained by Bera et al. (2022).
@@
We notice that our results have a non-negligible dependence on the model adopted to describe the $M_{\rm HI}-M_{\star}$ relation. In fact, the linear model tends to produce HIMF with higher normalization and larger values for $\Omega_{\rm HI}$ than the polynomial model. This is not surprising since, as discussed, the polynomial model steepens towards low $M_{\star}$ where a greater number of galaxies is found. In addition, a linear model extrapolates the $M_{\rm HI}-M_{\star}$ relation to arbitrarily high value of $M_{\rm HI}$. However, results at $z\sim 0$ show that massive galaxies feature a flattening in $M_{\rm HI}$ at the high-mass end of the $M_{\rm HI}-M_{\star}$ relation, due to a plethora of phenomena shaping galaxy evolution, mainly gas depletion via star formation. In this work, we find evidence for the polynomial model to better fit the data points from stacking. In particular, the best-fitting second-order polynomial model for the $M_{\rm HI}$ relation is found to have second-order coefficient $a<0$ at a significance $10\sigma$ in all the studied cases. While we regard this finding as non-conclusive given the limited amount of data points available for the fit, the trend suggested by the best-fitting polynomial fit qualitatively reproduces the one described by a double power law at $z\sim 0$ (Maddox et al., 2015; Parkash et al., 2018; Pan et al., 2023). In particular, the polynomial model with $a<0$ accounts for the high-mass flattening of the $M_{\rm HI}-M_{\star}$ scaling relation. For all these reasons we regard the polynomial model as the fiducial model for this work.
@@ 5.4 Comparison with simulations
All the other feedback models converge well within each other and are in good agreement with our results at $\log_{10}(M_{\rm HI}/{\rm M_{\odot}})\sim 9.7$, but tend to diverge towards high $M_{\rm HI}$ values. The noX model underpredicts our results, while the noAGN is in good agreement with them, and the noJet mode overpredicts the HIMF, respectively. This indicates that, as anticipated, the specific AGN feedback model is a major driver of the massive end of the HIMF. In particular, it is interesting to notice that the noAGN model reproduces very well the HIMF from this work. The finding suggests a scenario in which neglecting the feedback from supermassive black holes allows to better model the HIMF at the probed H I masses. This is consistent with the results from Gensior et al. (2023, 2024), where the authors considered tens of Milky-Way-like galaxies from the EMP-Pathfinder (Reina-Campos et al., 2022) and FIREbox (Feldmann et al., 2023) cosmological simulations and showed that they reproduce well the observed H I properties of nearby galaxies without the need of AGN feedback. Therefore, despite the large uncertainty associated to feedback models in current cosmological hydrodynamic simulations, we argue that a detailed study of these feedback mechanisms is crucial to properly understand which physical mechanism shapes the HIMF at the different H I masses.
@@ 6 Summary and conclusions
In this work we report a novel indirect semi-empirical estimation of the HIMF and of the H I cosmic density parameter $\Omega_{\rm HI}$ for star-forming galaxies at $z\sim 0.37$. In particular, we combine the most recent measurements of the stellar mass function of star-forming galaxies in the COSMOS field at $0.2<z<0.5$ (Weaver et al., 2023) and of the $M_{\rm HI}-M_{\star}$ scaling relation of star-forming galaxies at $0.22<z<0.49$ (median redshift $z\sim 0.37$) by Bianchetti et al. (2025) from the combined H I stacking analysis of two state-of-the-art deep H I interferometric surveys MIGHTEE (Jarvis et al., 2016) and CHILES (Fernández et al., 2013), both covering the COSMOS field. In particular, we adopt a Monte Carlo approach following the one pioneered by Bera et al. (2022) consisting of drawing samples of the stellar mass function, generate a synthetic stellar mass sample from each stellar mass function, and convert $M_{\star}$ to $M_{\rm HI}$ by assuming a $M_{\rm HI}-M_{\star}$ scaling relation with an adequate scatter. We compute the median and the standard deviation of the sample of HIMF, which defines our ‘measured’ HIMF. In addition, for each H I mass sample, we build the corresponding HIMF and fit a Schechter profile via a Markov Chain Monte Carlo procedure using the emcee affine-invariant sampler (Foreman-Mackey et al., 2013). This allows us to determine the full posterior distributions of the best-fitting Schechter parameters and hence, to determine not only the median values but also their associated uncertainties after having adequately propagated all the uncertainties in our forward modelling.
@@
  • we model the $M_{\rm HI}-M_{\star}$ scaling relation by fitting the data points from Bianchetti et al. (2025) using both a linear and a second-order polynomial model. In particular, the second-order coefficient $a$ is found to be negative, $a<0$. The resulting model resembles the well-known ‘curved trend’ observed in the $M_{\rm HI}-M_{\star}$ relation at $z=0$;
  • we find a global difference in the normalization of the HIMF from the $z\sim 0$ to $z\sim 1$. In particular, the resulting HIMF from this work has larger normalization than the one reported by ALFALFA at $z\sim 0$, and smaller normalization than the one reported by Chowdhury et al. (2024) at $z\sim 1$. We tentatively interpret this fact as as a signature of evolution of the HIMF, although it is hard to draw conclusions in this regard given the potential selection biases and systematics that could be affecting the comparison between our findings and the other literature results;
@@
  • the resulting values for $\Omega_{\rm HI}$ obtained by assuming the linear and the polynomial model for the scaling relation differ by $\sim 2.4\sigma$ in the combined stacking case.
@@ Appendix A Consistency of the HIMF from MIGHTEE and CHILES
Survey $\log_{10}(\tilde{M}/{\rm M}_{\odot})$ $\alpha$ $\log_{10}(\tilde{\Phi}/{\rm Mpc}^{-3})$ $\Omega_{\rm HI}\times 10^{-4}$
@@
MIGHTEE $10.13^{+0.08}_{-0.09}$ $1.38^{+0.11}_{-0.10}$ $2.24^{+0.13}_{-0.13}$ $8.16^{+0.59}_{-0.54}$
CHILES $10.01^{+0.16}_{-0.15}$ $1.31^{+0.23}_{-0.20}$ $2.18^{+0.21}_{-0.25}$ $6.65^{+0.99}_{-0.69}$
Combined $10.16^{+0.08}_{-0.08}$ $1.48^{+0.10}_{-0.09}$ $2.41^{+0.12}_{-0.13}$ $7.02^{+0.59}_{-0.52}$
@@
MIGHTEE $10.01^{+0.10}_{-0.09}$ $1.27^{+0.17}_{-0.16}$ $1.96^{+0.14}_{-0.16}$ $1.04^{+0.94}_{-0.79}$
CHILES $9.95^{+0.17}_{-0.16}$ $1.30^{+0.28}_{-0.25}$ $2.01^{+0.22}_{-0.27}$ $8.49^{+1.70}_{-1.16}$
Combined $10.13^{+0.09}_{-0.09}$ $1.56^{+0.14}_{-0.13}$ $2.31^{+0.15}_{-0.17}$ $9.98^{+0.11}_{-0.18}$
@@
The best-fitting values for the parameters of the Schechter model are always in agreement within $1\sigma$ in the combined, MIGHTEE, and CHILES cases. When comparing the results from the linear and the polynomial model, we report a systematic shift of $\log_{10}(\tilde{M}/{\rm M}_{\odot})$ towards larger values and of $\alpha$ and $\log_{10}(\tilde{\Phi}/{\rm Mpc}^{-3})$ towards smaller values when passing from the linear to the polynomial model. While these shifts are not statistically significant and are often $1\sigma$, we observe them systematically for all the three cases. In addition, we show in Appendix F the posterior distributions for the different cases. Figs. 10 and 11 show the resulting posterior distributions for the model parameters of the Schechter profile for the linear and the polynomial model, respectively. The distributions have qualitatively similar shapes, and the ones for CHILES are manifestly broader than the posteriors of the MIGHTEE and the combined cases, as expected from the smaller statistics. We will have a closer look to the effect of the sample size in Appendix E. Figs. 12, 13, and 14 display instead a comparison between the posterior distributions from the linear (blue solid) and the polynomial model (orange dashed), for the combined, MIGHTEE, and CHILES cases, respectively. Here, we clearly appreciate the systematic shift of the posteriors when passing from the linear to the polynomial model.
@@ Appendix B The impact of scatter
[…]
@@
MIGHTEE $10.13^{+0.08}_{-0.09}$ $1.38^{+0.11}_{-0.10}$ $2.24^{+0.13}_{-0.13}$ $8.16^{+0.59}_{-0.54}$
CHILES $10.01^{+0.16}_{-0.15}$ $1.31^{+0.23}_{-0.20}$ $2.18^{+0.21}_{-0.25}$ $6.65^{+0.99}_{-0.69}$
Combined $10.16^{+0.08}_{-0.08}$ $1.48^{+0.10}_{-0.09}$ $2.41^{+0.12}_{-0.13}$ $7.02^{+0.59}_{-0.52}$
@@
MIGHTEE $10.11^{+0.09}_{-0.10}$ $1.37^{+0.11}_{-0.10}$ $2.23^{+0.12}_{-0.14}$ $8.09^{+0.60}_{-0.54}$
CHILES $10.02^{+0.16}_{-0.16}$ $1.33^{+0.22}_{-0.19}$ $2.20^{+0.21}_{-0.24}$ $6.67^{+0.98}_{-0.80}$
Combined $10.15^{+0.08}_{-0.08}$ $1.48^{+0.10}_{-0.09}$ $2.41^{+0.13}_{-0.13}$ $6.99^{+0.61}_{-0.51}$
@@ Appendix C The impact of the $M_{\rm HI}$ distribution asymmetry
Figure 7: Results of testing the impact of an asymmetric scatter for the $M_{\rm HI}-M_{\star}$ relation. Top: $M_{\rm HI}$ distributions (to which we have subtracted the median) in different stellar mass bins: $8.9<\log_{10}(M_{\star}/{\rm M_{\odot}})<9.1$ (green solid), $9.4<\log_{10}(M_{\star}/{\rm M_{\odot}})<9.6$ (orange dashed), $9.9<\log_{10}(M_{\star}/{\rm M_{\odot}})<10.1$ (blue dotted-dashed), $10.4<\log_{10}(M_{\star}/{\rm M_{\odot}})<10.5$ (purple dotted). We display as a gray dashed a lognormal random distribution obtained by assuming $\sigma=0.45$, which is found to approximate well all the $M_{\rm HI}$ distributions.
@@ Appendix D Completeness effects on the HIMF
Fig. 8 shows the resulting HIMF from assuming the combined scaling relation and the linear (yellow stars) and the polynomial (blue squares) models. The HIMF presents evidence of incompleteness at $\log_{10}(M_{\rm HI}/{\rm M}_{\odot})\lesssim 9$. As stated in Sect. 4.3, we conservatively fit the Schechter model only at $\log_{10}(M_{\rm HI}/{\rm M}_{\odot})>9.2$ for the polynomial model and at $\log_{10}(M_{\rm HI}/{\rm M}_{\odot})>9.4$ for the linear model. The $M_{\rm HI}$ lower limits for the fit, $\log_{10}(M_{\rm HI}/{\rm M}_{\odot})=9.2$ and $\log_{10}(M_{\rm HI}/{\rm M}_{\odot})=9.4$, are marked by a vertical black dashed line and a vertical black dotted line, respectively. We display the best-fitting Schechter models obtained adopting the fitting strategy outlined above as solid yellow (linear model) and solid blue (polynomial model lines.
Figure 8: Completeness effects on the HIMF. We display the resulting HIMF obtained by assuming the linear (yellow stars) and the polynomial (blue squares) models for the $M_{\rm HI}-M_{\star}$ relation. The vertical gray dashed and dotted line stand for the conservative completeness limits $\log_{10}(M_{\rm HI}/{\rm M}_{\odot})=9.2$ and $\log_{10}(M_{\rm HI}/{\rm M}_{\odot})=9.4$ for the polynomial and linear cases, respectively, and represent the lower limits used for the fit of the Schechter models to the data. We show the best-fitting Schechter models obtained adopting the fitting strategy outlined above as solid yellow (linear model) and solid blue (polynomial model lines. We also display the HIMF from ALFALFA at $z\sim 0$ (green dashed) for comparison.